the magnetic flux through the loop increases according to the relation ΦB = 8.0t2 + 5.9t, where ΦB is in milliwebers and t is in seconds. (a) What is the magnitude of the emf induced in the loop when t = 2.9 s? (b) Is the direction of the current through R to the right or left?
Answer #1

(as Magnitudes of Induced cemf = doo E = + (8+2+5.9+) d- = 16t + 5.9 V at t=2.95 E = (1682.9+5.g)v=5203 v (6) The direction o